OpenStudy (anonymous):

Emma is going on vacation this summer and is trying to choose between two different travel plans. The first plan costs $450 for 3 days at a hotel and 2 days at King's Island. The second plan offers 5 days at the same hotel and 3 days at King's Island for $700. The cost of at the hotel and King's Island is the same under both plans. How much does a 1-day trip to King's Island cost?

OpenStudy (anonymous):

@fallonbanks

OpenStudy (anonymous):

@One098

OpenStudy (one098):

Do you have any idea as to what the answer might be?

OpenStudy (anonymous):

I know how to find the answer, I just don't know how to set up the equations.

OpenStudy (one098):

Alright, give me a sec.

OpenStudy (one098):

Well you can make the first one something like: 450=3h+2k h=cost per night at the hotel k=cost per night at King's Island

OpenStudy (one098):

or it should be more like the cost per day because that's what it implies.. "cost per day."

OpenStudy (anonymous):

So the second one would be 700= 5h + 3k?

OpenStudy (one098):

Exactly what I was about to say! So, yes!

OpenStudy (anonymous):

Alright, I got it, thanks!

OpenStudy (one098):

My pleasure, anytime!

OpenStudy (anonymous):

I don't got it, what's next?

OpenStudy (one098):

Lol. Alright, well you first want to solve for h, okay? So isolate the h.

OpenStudy (one098):

Wait, no. Move the h to the left side of the equal sign.

OpenStudy (anonymous):

Still lost?

OpenStudy (one098):

Hm alright then. \[450=3h+2k\] Divide both sides by 3h: \[\frac{ 450 }{ 3h }=\frac{ 3h }{ 3h }+2k\] The 3h/3h cancel out.

OpenStudy (anonymous):

Brilliant!

OpenStudy (one098):

Do you know the next step?

OpenStudy (anonymous):

No, sorry.

OpenStudy (anonymous):

brb

OpenStudy (one098):

Now you have: \[\frac{ 450 }{ 3h }=2k\]

OpenStudy (anonymous):

That's the equation i use to find the answer??

OpenStudy (one098):

Yes and you must do the same to find the next equation, however I think there is another way, its simpler, to find the answers..

OpenStudy (anonymous):

Which is what?

OpenStudy (one098):

\[450=3h+2k\] Instead, subtract 2k from both sides.

OpenStudy (anonymous):

2k-450=3h?

OpenStudy (one098):

Well you see, the 450 isn't a negative, so it must be positive. \[450-2k=3h\]